OpenStudy (anonymous):

find the zeros y=x^2+9x+8

OpenStudy (openstudier):

We can factor this, do you know to factor?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

k

OpenStudy (anonymous):

y=x^2+9x+8 y=(x+8)(x+1) when y=0; (x+8)(x+1)=0 x=-8,-1 :. y=0, when x=-8,-1.